OpenStudy (anonymous):

How do you find the dot product of:

OpenStudy (anonymous):

a*b a=6i+5j b=-5i+4j

OpenStudy (anonymous):

u have to simplify

OpenStudy (anonymous):

I did i tried the one where you expand it and I got 24+25 and that's49 it wasn't one of my choices

OpenStudy (anonymous):

Then i did 6*-5,5*4 but I know that's wrong bc ive done it before like that and i got it wrong

OpenStudy (anonymous):

ooo ok thati dont kow i just kow you culd simplyfi sorry

OpenStudy (anonymous):

how would you do it though?

OpenStudy (anonymous):

i dont kow im going to tell some body else to help u in a second ok

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

6*-5+5*4=-30+20=-10

OpenStudy (dumbcow):

hmm no you are right .... its -10

OpenStudy (anonymous):

<(x1,x2),(y1,y2)> = x1*y1+x2*y2

OpenStudy (anonymous):

I was close i just had to minus it thank you :)